OpenStudy (anonymous):

Solve for x: √2x+5-1=x

ganeshie8 (ganeshie8):

\(\large \sqrt{2x+5} - 1 = x\)

ganeshie8 (ganeshie8):

like that ?

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

add 1 both sides, square both sides

ganeshie8 (ganeshie8):

\(\large \sqrt{2x+5} - 1 = x \) +1 +1

OpenStudy (nurali):

√2x+5-1=x √2x+5=x+1 squaring both side 2x+5=(x+1)^2 2x+5=x^2+2x+1 x^2+2x-2x-5+1=0 x^2-4=0 x^2=4 x=+-sqrt(4) x=+-2 x=2 or x=-2

OpenStudy (anonymous):

omg, thank you so much for showing a complete step by step! i can put it in my notebook for notes of next time

OpenStudy (nurali):

Anytime.